OpenStudy (susanna):

What is the equation for the inverse of y = 1/2 Sin x? Is the answer y=2 Arcsin x?

OpenStudy (mathstudent55):

Switch x and y in the original equation, and solve for y.

OpenStudy (mathstudent55):

Start with the first step and show it. Write the original equation and switch the variables.

OpenStudy (michele_laino):

hint: we can write this: \[\Large y = \frac{1}{2}\sin x \Rightarrow \sin x = 2y\]

OpenStudy (susanna):

y=2 Arcsin x ?

OpenStudy (mathstudent55):

To answer your question, the answer is no, that is not the answer. How about paying attention to what I am trying to show you. I am guiding you through the process of finding an inverse function. Start with the original function. Now switch the variables, what do you get?

OpenStudy (susanna):

x=1/2 Sin y

OpenStudy (susanna):

Arcsin x=1/2 y

OpenStudy (mathstudent55):

Great. (The first of the two lines above.) Now you need to solve for y. Switch sides to have the y on the left side. Then multiply both sides by 2.

OpenStudy (susanna):

I got 2 Arcsin x=y

OpenStudy (mathstudent55):

One step at a time. This is good. \(x = \dfrac{1}{2} \sin y\) Now switch sides. Then multiply both sides by 2.

OpenStudy (susanna):

switch sides?

OpenStudy (mathstudent55):

Switch the left and right sides of the equation so you have the y variable on the left side.

OpenStudy (mathstudent55):

\(x = \dfrac{1}{2} \sin y\) \(\dfrac{1}{2} \sin y = x\)

OpenStudy (mathstudent55):

Since you have 1/2 multiplying sin y, and you want to solve for sin y then for y, multiply both sides by 2.

OpenStudy (susanna):

sin y=2x

OpenStudy (mathstudent55):

Great. Now finally, you take the inverse sin, or the arcsin.

OpenStudy (susanna):

y=arcsin 2x

OpenStudy (mathstudent55):

Exactly. Now you got it right. Good job!

OpenStudy (susanna):

thanx

OpenStudy (susanna):

I like the way you explain it, instead of giving a straight answer.

OpenStudy (susanna):

Now I know how to solve these kinds of problems.

OpenStudy (mathstudent55):

The process is always the same to find an inverse function: Step 1. Write the original equation. If the function is given in function notation, \(f(x) = ...\) , then replace f(x) with y. Step 2. Switch x and y. Step 3. Solve for y. Step 4. If you need function notation, replace y with \(f^{-1}(x) \)

OpenStudy (mathstudent55):

You're welcome.